# Heavy neutral bosons and dark matter in the 3-3-1 model with axionlike particle

### v1

@@ III.1 Couplings and Analytical form related to LFV processes
$+$ $\left(y_{\nu}^{D}\right)_{ab}\left(\overline{\nu}_{aL}\eta_{1}^{0}+\overline{l}_{aL}\eta_{2}^{-}+\overline{\nu_{R}^{c}}_{aL}\eta_{3}^{0}\right)N_{bR}+\mbox{H.c.},$
Accordingly, the mass mixing matrices of the leptons are given:
@@
$g_{ab}=\frac{\sqrt{2}m_{l_{a}}}{v_{\rho}}\delta_{ab},\hskip 8.5359pty^{D}_{\nu}=\frac{\sqrt{2}M^{D}_{\nu}}{v_{\eta}},\hskip 8.5359pty_{N}=\frac{M_{N}}{\sqrt{2}v_{\phi}},$ (86)
@@
$\lambda_{h_{1}H^{+}_{2}H^{-}_{2}}$ $=s_{\alpha_{2}}\left(c_{\theta_{1}}s_{\theta_{1}}v_{\chi}\lambda_{8}+c_{\alpha}v\left(2\lambda_{2}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\left(\lambda_{5}+\lambda_{8}\right)\right)\right)+c_{\alpha_{2}}\left(s_{\alpha}v\left(\lambda_{6}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\lambda_{5}\right)-c_{\theta_{1}}s_{\theta_{1}}v_{\phi}\lambda_{\phi}\right),$ (91)
@@
Therefore, Eq.(96) is rewritten as Eq.(97, where $g_{HGG}$ is given in table 2 and $g_{Hff}$ as in table 4. While $g^{SM}_{hGG}$ and $g^{SM}_{hff}$ are the corresponding couplings of the SM Higgs boson.
@@ III.2 Numerical results
Moreover, the interaction constants are chosen as dependent parameters based on Eqs.(57,75), the rest will be treated as free parameters and must be in the range of values that ensure the limits of perturbation theory ($\lambda_{i}\leq 4\phi$) . These dependent parameters are listed as in Eq. (III.2).
@@
By using the neutrino oscillation data as shown in Ref.Zyla:2020zbs ; Patrignani:2016xqp : $s^{2}_{12}=0.32,\;s^{2}_{23}=0.551,\;s^{2}_{13}=0.0216,\\ \Delta m^{2}_{21}=7.55\times 10^{-5}\;\mathrm{eV^{2}},\hskip 8.5359pt\Delta m^{2}_{32}=-2.50\times 10^{-3}\;\mathrm{eV^{2}}$, and choosing the Dirac and Majorana phases of the $U_{PMNS}$ in their simplest form (Refs.Quyet:2023dlo ; Hung:2022vqx ; Hue:2021zyw ), we can get the right part of Eq.(102) in the known form. Furthermore, $M_{\nu}^{D}$ must be antisymmetric so it contains only six unknowns. Combined with $M_{N}$ in diagonal form, we get three more unknowns. There are a total of nine unknowns on the right part, so Eq.(102) can be solved. That means we have successfully parametrized $U^{\nu}$.
@@
Based on the analytical form of the amplitudes given in App.B, we investigate the values of $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$, which are given in Fig.3. We can see that, most of the values obtained by $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$, in left panel, and $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$, in right panel, are within the current experimental limits ($\leq 10^{-3}$).
@@
$\sigma^{h}\left(pp\rightarrow H\right)=\mathcal{L}_{q\overline{q}}\left(m_{H}^{2}\right)\hat{\sigma}\left(\overline{q}_{i}q_{i}\rightarrow H\right)$ (103)
@@
Figure 4: Plots of $\sigma\left(pp\rightarrow h_{2}\right)\times\mathrm{Br}\left(h_{2}\rightarrow\mu\tau\right)$ (magenta line) depend on $M_{H}$, the green and cyan lines depict the CMS’s observations for the combined decay modes of $H\rightarrow\mu\tau$ (based on Ref. CMS:2019pex ) for the low and high mass ranges, respectively..
@@
IV $Z^{\prime}$ boson and dark matter
@@
$\mathcal{L}^{Z^{\prime}}_{NC}\supset\overline{f}\gamma^{\mu}\left(g_{V}^{Z^{\prime}}(f)+g_{A}^{Z^{\prime}}(f)\gamma^{5}\right)fZ^{\prime}_{\mu}$ (107)
where $g_{V}^{Z^{\prime}}$ and $g_{A}^{Z^{\prime}}$ are given in Tab.5,
@@
$Z^{\prime}H^{+}_{2}Y^{-}$ $\frac{g^{2}\left(c_{\alpha}c_{\theta_{1}}\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)-24s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)v+\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)+12s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)s_{\theta_{1}}v_{\chi}\right)}{12\sqrt{4c_{W}^{2}-1}}$
@@
Because $Z$ and $Z^{\prime}$ do not mix as shown in Eq.(II.2), $\Gamma(Z^{\prime}\rightarrow W^{+}W^{-})$ is canceled out according to Ref.Perez:2004jc . Furthermore, $m_{X}>m_{Z^{\prime}}/2$ means $\Gamma(Z^{\prime}\rightarrow X^{*}X,\penalty 10000\ X\equiv Y^{\pm},X^{0*},X^{0}$ also do not contribute to the total width of $Z^{\prime}$. The remaining components are in the form as follows Ninh:2005su ; Perez:2004jc :
@@
Next, we consider the DM candidate in the model. By introducing new quantum numbers, the relation to the usual B (L) is cl ; Long:2024rdq :
$B=\mathcal{B}\text{\bf I},\penalty 10000\ L=-\frac{4}{\sqrt{3}}T_{8}+\mathcal{L}\text{\bf I}$ (115)
We add the symmetry group $U(1)_{\mathcal{N}}\otimes\mathcal{Z}_{2}^{\prime}$ to the model, with $\mathcal{N}=3\left(\mathcal{B}-\mathcal{L}\right)$ and $\mathcal{Z}_{2}^{\prime}$ being the discrete group structured as,
$\mathcal{Z}_{2}^{\prime}=\left\{1,g\right\},\text{with}\penalty 10000\ g=(-1)^{\mathcal{N}+2s}$ (116)
where $p_{s}=(-1)^{2s}$ is spin parity.
Thus, $\mathcal{Z}_{2}^{\prime}$ posses two one-dimensional irreducible representations, $\underline{1}$ according to $g=1$ and $\underline{1^{\prime}}$ according to g=1g=-1. Under the transformation of $\mathcal{Z}_{2}^{\prime}$, the particles are classified into even and odd according to their charge, all SM particles have even charges, denoted ++. Other fields are also determined by charge of $\mathcal{Z}_{2}^{\prime}$ through $\mathcal{B}$ and $\mathcal{L}$, the results are given as in table 7.
Multiplets $Q_{3L}$ $Q_{nL}$ $u_{aR}$ $d_{aR}$ $U_{3R}$ $D_{nR}$ $\psi_{aL}$ $l_{aR}$ $\eta$ $\rho$ $\chi$ $\phi$ $N_{aR}$
$\mathcal{B}$ $\frac{1}{3}$ $\frac{1}{3}$ $\frac{1}{3}$ $\frac{1}{3}$ $\frac{1}{3}$ $\frac{1}{3}$ 0 0 0 0 0 0 0
$\mathcal{L}$ $\frac{2}{3}$ $\frac{2}{3}$ 0 0 2-2 22 $\frac{1}{3}$ 11 $\frac{2}{3}$ $\frac{2}{3}$ $\frac{4}{3}$ 0 0
$\mathcal{N}$ 1-1 33 11 11 5-5 77 11 3-3 2-2 2-2 44 0 0
$\mathcal{Z}_{2}^{\prime}$ ++ ++ ++ ++ ++ ++ ++ ++ ++ ++ ++ ++ -
Table 7: The $U(1)_{\mathcal{N}}\otimes\mathcal{Z}_{2}^{\prime}$ charges of fields.
It should be emphasized that the appearance of the $U(1)_{\mathcal{N}}\otimes\mathcal{Z}_{2}^{\prime}$ group does not affect the couplings in Tab.(5,6) nor the other interactions related to Fig.5. This is simple because almost fields are even of $\mathcal{Z}_{2}^{\prime}$. The difference lies in $N_{aR}$. To clarify this, we rely on the chiral anomalies elimination condition of the model. While the triangular anomalies are eliminated, two non-trivial ones related to $U(1)_{\mathcal{N}}$ which are
$\left[Gravity\right]^{2}U(1)_{\mathcal{N}}\sim\sum_{a}\mathcal{N}_{N_{aR}}=0,\penalty 10000\ U(1)_{\mathcal{N}}^{3}\sim\sum_{a}\mathcal{N}_{N_{aR}}^{3}=0$ (117)
The only way to satisfy Eq.(117), that $N_{aR}$ get zero-charge of $U(1)_{\mathcal{N}}$, it cause we have three odd- $\mathcal{Z}_{2}^{\prime}$ fields. Furthermore, $\mathcal{Z}_{2}^{\prime}$ is a residual symmetry after spontaneous breaking, which is a crucial thing in proposing $N_{aR}$ as dark matter candidates.
Among the odd- $\mathcal{Z}_{2}^{\prime}$ particles, we assume $N_{1R}$ is the lightest. So, $N_{1R}$ is considered dark matter and its pair annihilates in to the SM particles in the early universe. Because $N_{1R}$ only direct interacts with $\Phi$ according to Eq.(25), but $\Phi$ hardly mixes with other fields Eq.(54), the contribution to the annihilation of dark matter particles mainly comes from the s-channel exchange diagrams mediated by $\Phi$. The diagram is shown in Figure 6.
@@
where $x_{f}\simeq 20.25$ Bertone:2004pz , and $g_{\Phi ZZ}$, $g_{\Phi h_{1}h_{1}}$ are determined from Eq.(26) and Eq.(54), respectively.
$\Omega_{DM}h^{2}\simeq\frac{0.1\penalty 10000\ \mathrm{pb}}{\langle\sigma v_{rel}\rangle_{N_{1R}^{C}N_{1R}\rightarrow all}}$ (119)
@@
V Conclusions
In the 3-3-1 model with axionlike particles (ALP), we show the region of the parameter space where both cLFV and LFVHDs are satisfied experimental limits. Based on gluon-gluon fusion at the LHC, we investigate the signals of cross-sections in the allowed parameter space, and provide predictions of $m_{h_{2}}\geq 600\penalty 10000\ \mathrm{GeV}$. Based on the search for high-mass dilepton resonances at ATLAS and CMS, the mass of $Z^{\prime}$-boson is predicted as $m_{Z^{\prime}}\geq 5.1\penalty 10000\ \mathrm{TeV}$. We introduced the symmetry group $U(1)_{\mathcal{N}}\otimes\mathcal{Z}_{2}^{\prime}$ into the model to indicate dark matter candidates. Investigating the relic density of dark matter within experimentally permissible limits, we also established a relationship between the mass of dark matter and the breaking scale of axion.
@@ Acknowledgments
H.T. Hung thanks to A.V. Bednyakov for the helpful comments to improving this article.

### v2

@@ III.1 Couplings and Analytical form related to LFV processes
$+$ $\left(y_{\nu}^{D}\right)_{ab}\left(\overline{\nu}_{aL}\eta_{1}^{0}+\overline{l}_{aL}\eta_{2}^{-}+\overline{\nu_{R}^{c}}_{aL}\eta_{3}^{0}\right)N_{bR}+\mbox{H.c.}$
The mass mixing matrices of the leptons are given:
@@
$g_{ab}=\frac{\sqrt{2}m_{l_{a}}}{v_{\rho}}\delta_{ab},\hskip 8.5359pty^{D}_{\nu}=\frac{\sqrt{2}M^{D}_{\nu}}{v_{\eta}},\hskip 8.5359pty_{N}=\frac{M_{N}}{\sqrt{2}v_{\phi}}.$ (87)
@@
$\lambda_{h_{1}H^{+}_{2}H^{-}_{2}}$ $=s_{\alpha_{2}}\left(c_{\theta_{1}}s_{\theta_{1}}v_{\chi}\lambda_{8}+c_{\alpha}v\left(2\lambda_{2}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\left(\lambda_{5}+\lambda_{8}\right)\right)\right)+c_{\alpha_{2}}\left(s_{\alpha}v\left(\lambda_{6}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\lambda_{5}\right)-c_{\theta_{1}}s_{\theta_{1}}v_{\phi}\lambda_{\phi}\right).$ (92)
@@
Therefore, Eq.(97) is rewritten as Eq.(98), where $g_{HGG}$ is given in table 2 and $g_{Hff}$ as in table 4. While $g^{SM}_{hGG}$ and $g^{SM}_{hff}$ are the corresponding couplings of the SM Higgs boson.
@@ III.2 Numerical results
Moreover, the interaction constants are chosen as dependent parameters based on Eqs.(58,76), the rest will be treated as free parameters and must be in the range of values that ensure the limits of perturbation theory ($\lambda_{i}\leq 4\pi$) . These dependent parameters are listed as in Eq. (III.2).
@@
By using the neutrino oscillation data as shown in Ref.Zyla:2020zbs ; Patrignani:2016xqp ; Tanabashi:2018oca ; ParticleDataGroup:2024cfk ; JUNO:2021vlw : $s^{2}_{12}=0.307,\;s^{2}_{23}=0.558,\;s^{2}_{13}=0.0219,\\ \Delta m^{2}_{21}=7.53\times 10^{-5}\;\mathrm{eV^{2}},\hskip 8.5359pt\Delta m^{2}_{32}=2.455\times 10^{-3}\;\mathrm{eV^{2}}$, and choosing the Dirac and Majorana phases of the $U_{PMNS}$ in their simplest form (Refs.Quyet:2023dlo ; Hung:2022vqx ; Hue:2021zyw ), we can get the right part of Eq.(103) in the known form. Furthermore, $M_{\nu}^{D}$ must be antisymmetric so it contains only six unknowns. Combined with $M_{N}$ in diagonal form, we get three more unknowns. There are a total of nine unknowns on the right part, so Eq.(103) can be solved. That means we have successfully parametrized $U^{\nu}$.
@@
Based on the analytical form of the amplitudes given in appendix B, we investigate the values of $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$, which are given in Fig.3. We can see that, most of the values obtained by $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$, in left panel, and $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$, in right panel, are within the current experimental limits ($\leq 10^{-3}$).
@@
$\sigma^{h}\left(pp\rightarrow H\right)=\mathcal{L}_{q\overline{q}}\left(m_{H}^{2}\right)\hat{\sigma}\left(\overline{q}_{i}q_{i}\rightarrow H\right).$ (104)
@@
Figure 4: Plots of $\sigma\left(pp\rightarrow h_{2}\right)\times\mathrm{Br}\left(h_{2}\rightarrow\mu\tau\right)$ (magenta line) depend on $M_{H}$, the green and cyan lines depict the CMS’s observations for the combined decay modes of $H\rightarrow\mu\tau$ (based on Ref. CMS:2019pex ) for the low and high mass ranges, respectively.
@@
IV $Z^{\prime}$ boson
@@
$\mathcal{L}^{Z^{\prime}}_{NC}\supset\overline{f}\gamma^{\mu}\left(g_{V}^{Z^{\prime}}(f)+g_{A}^{Z^{\prime}}(f)\gamma^{5}\right)fZ^{\prime}_{\mu},$ (108)
where $g_{V}^{Z^{\prime}}$ and $g_{A}^{Z^{\prime}}$ are given in table 5.
@@
$Z^{\prime}H^{+}_{2}V^{-}$ $\frac{g^{2}\left(c_{\alpha}c_{\theta_{1}}\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)-24s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)v+\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)+12s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)s_{\theta_{1}}v_{\chi}\right)}{12\sqrt{4c_{W}^{2}-1}}$
@@
Because $Z$ and $Z^{\prime}$ do not mix as shown in Eq.(II.2), $\Gamma(Z^{\prime}\rightarrow W^{+}W^{-})$ is canceled out according to Ref.Perez:2004jc . Furthermore, $m_{X}>m_{Z^{\prime}}/2$ means $\Gamma(Z^{\prime}\rightarrow X^{*}X,\penalty 10000\ X\equiv V^{\pm},X^{0*},X^{0}$ also do not contribute to the total width of $Z^{\prime}$. The remaining components are in the form as follows Ninh:2005su ; cl ; Perez:2004jc :
@@
V Dark matter
From the symmetry breaking scheme in (II.1), we see that $Z_{2}$ is the residual symmetry of the model, so odd-$Z_{2}$ particles can be considered as dark matter candidates. These include: $\eta,\,\rho,\,u_{aR},\,d_{nR},\,l_{aR},\,N_{aR}$ according to (4). But $\eta$ and $\rho$ directly break electroweak scale, and $u_{aR},\,d_{nR},\,l_{aR}$ are components that create ordinary matter, so only $N_{aR}$ remains as the best choice to consider as dark matter candidates in the model.
Among $N_{aR}$, we assume $N_{1R}$ is the lightest. So, $N_{1R}$ is considered dark matter and its pair annihilates in to the SM particles in the early universe. Because $N_{1R}$ only direct interacts with $\Phi$ according to Eq.(25), and $\Phi$ can couple with SM particles as Eq.(55), the contribution to the annihilation of dark matter particles mainly comes from the s-channel exchange diagrams mediated by $\Phi$. The diagram is shown in Figure 6.
@@ V Dark matter
where $x_{f}\simeq 20.25$ Bertone:2004pz and $g_{\Phi N_{1R}^{C}N_{1R}}=\left(y_{N}\right)_{11}$, while on the contrary $g_{\Phi ZZ}=0$, $g_{\Phi h_{1}h_{1}}=\frac{v_{\phi}}{2\sqrt{2}}\left(\lambda_{12}s_{\alpha_{2}}^{2}+\lambda_{13}c_{\alpha_{2}}^{2}\right)$ are determined from Eq.(27) and Eq.(55), respectively.
$\Omega_{DM}h^{2}\simeq\frac{0.1\penalty 10000\ \mathrm{pb}}{\langle\sigma v_{rel}\rangle_{N_{1R}^{C}N_{1R}\rightarrow all}}.$ (117)
@@
VI Conclusions
In the 3-3-1 model with axionlike particles (ALP), we show the region of the parameter space where both cLFV and LFVHDs are satisfied experimental limits. These are allowed parameter spaces where $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ has a value about of $10^{-6}$ and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$ reaches a value approximately of $10^{-5}$. Based on gluon-gluon fusion at the LHC, we investigate the signals of cross-sections in the allowed parameter space, and provide predictions of $m_{h_{2}}\geq 600\penalty 10000\ \mathrm{GeV}$. Based on the search for high-mass dilepton resonances at ATLAS and CMS, we take the numerical result of $\sigma\left(pp\rightarrow Z^{\prime}\rightarrow\overline{l}l\right)$ and show that the mass of $Z^{\prime}$-boson is $m_{Z^{\prime}}\geq 5.1\penalty 10000\ \mathrm{TeV}$. The hypothesis of the existence of residual symmetry $Z_{2}$ after spontaneous symmetry breaking stages allows that only $N_{aR}$ are dark matter candidates among the odd -$Z_{2}$ particles. Investigating the relic density of dark matter within experimentally permissible limits, we also established a relationship between the mass of dark matter and the breaking scale of axion.
@@ Acknowledgments
H.T. Hung would like to thank A.V. Bednyakov for the helpful comments to improving this article.